OpenStudy (anonymous):

find the cost of painting the sides of a tent measuring 4m in diameter and 5m in slant height at P25 per square meter.

OpenStudy (anonymous):

pls answer this

Directrix (directrix):

Lateral Area of Right Cone = pi * r * l where r is the radius of the circular base and l is the slant height. LA = pi* 2 * 5 = 10 pi square meters. *If P25* in the posted question means $25 per square meter, then the cost of the painting is 10 pi * 25 = $785.40 @ABBY20

OpenStudy (anonymous):

if i use 3.14 in pi will it be 78.5 dollars?

Directrix (directrix):

No. I used more digits but 3.14 is enough. Perhaps, I made an error. Question for you: What does P25 mean? I read it as $25. @Abby20

OpenStudy (anonymous):

yes

Directrix (directrix):

Did you multiply these numbers: 3.14 * 2 * 5 * $25? 3.14 * 2 * 5 * $25 = 785 U.S. dollars @ABBY20

OpenStudy (anonymous):

my equation is like this LA= pi*r*h LA=3.14*2*5 LA= 6.28*5 LA= 3.14sq.m. 25*3.14 = 78.5 dollars